OpenStudy (anonymous):

*Please Someone Help!* maximize P=1x+0.8y=0.7z Subject to 30x+20y+50z(less than or equalto)<40 5x+15y+17z(less than or equal to)<8 x+y+z(greater than or equal to)>3.5

OpenStudy (anonymous):

\[P=1x+0.8y+0.7z\] Is The First One, I Messed Up On The First Post

OpenStudy (anonymous):

This is a linear programming problem, and the best way to solve it is by using the simplex method. Do you know the method?

OpenStudy (anonymous):

great linear algebra question

OpenStudy (anonymous):

You could also draw it. P is a plane and the other 3 ones are 3-dimensional spaces, and you just need to find the intersection :)

OpenStudy (anonymous):

Yes, That is when I have to add slack Variables. But that's as far as I can get

OpenStudy (anonymous):

Good. You need to add only slack variables to the first two constrains since it's a maximization problem and they're less-than-or-equal constrains. But you need to add also an artificial variable to the last constrain.

OpenStudy (anonymous):

If you're not familiar with the term "artificial variable", then it's a variable you add in order to get a feasible solution to the objective function.

OpenStudy (anonymous):

30x + 20y + 50z + 1 + 0 + 0 + 0 = 40 5x + 15y + 17z +0 + 1 +0 + 0 =8 1x + 1y + 1z + 0 + 0 - 1 + 0 = 3.5 -1x - 0.8y - 0.7z + 0 + 0 + 0 + 1 = 0 Correct?

OpenStudy (anonymous):

Almost! \(30x + 20y + 50z + s_1 + 0s_2 + 0s_3 + 0a_1 = 40\) \(5x + 15y + 17z +0s_1 + s_2 +0s_3 + 0a_1 =8\) \(1x + 1y + 1z + 0 + 0 - 1s_3 + a_1 = 3.5\)

OpenStudy (anonymous):

Where \(s_i\) represents the slack variables, and \(a_1\) is an artificial variable. The objective function will then be \(P=x+0.8y+0.7z-Ma_1\), where M is a very large positive number.

OpenStudy (anonymous):

We never did the artificial numbers :-/

OpenStudy (anonymous):

You can watch this lecture for better understanding http://www.youtube.com/watch?v=wdGiekwXM2w&feature=relmfu

OpenStudy (anonymous):

Our professor completely skipped that

OpenStudy (anonymous):

Did you study the dual simplex method?

OpenStudy (anonymous):

You can use the dual simplex method to solve it if you know the method.

OpenStudy (anonymous):

Yes, we did

OpenStudy (anonymous):

You can use it, although I think it would be a bit lengthy.

OpenStudy (anonymous):

You need to multiply the last constrain by \(-1\), and then add the slack variables.

OpenStudy (anonymous):

Is the dual simplex method not only for minimization problems?

OpenStudy (anonymous):

Not really! You can actually always convert a maximization problem to a minimization problem by multiplying the objective function by \(-1\). I still think using the artificial variable is better here. It would take you less than 30 minutes to learn it.

OpenStudy (anonymous):

Let me ask you a question, are all the decision variables positive?

OpenStudy (anonymous):

I think they are

OpenStudy (anonymous):

@dmihail: I just saw your comment. Using graphical method with a three variables LP problem is difficult. It's probably the best method if you have two variables.

OpenStudy (anonymous):

Im attempting to watch the video you posted. Sure wish i had this video before

OpenStudy (anonymous):

Good luck. It won't be difficult to use artificial variables if you're already familiar with the method. :D

OpenStudy (anonymous):

If you worked that problem out, would it even work? last time i had a problem like this it was a trick question that had no solution

OpenStudy (anonymous):

What do you mean, would it even work?

OpenStudy (anonymous):

correct. Would there be an answer for x, y and z?

OpenStudy (anonymous):

I believe so!

OpenStudy (anonymous):

Awesome Thanks. Will chat back if the video confuses me LoL

OpenStudy (anonymous):

OK, if I'm still here :D

OpenStudy (anonymous):

How would the problem look written out in a matrix form?

OpenStudy (anonymous):

I'm not used to the matrix form, sorry!

OpenStudy (anonymous):

Could you break it down using the simplex method. Im going crazy trying to figure it out. I have been working on it for almost a week and cannot figure it out

OpenStudy (anonymous):

I'm busy right now. I'll write it down when I get time. When do you want the solution?

OpenStudy (anonymous):

You can email me at any time today... vdooley777@yahoo.com I appreciate all your help

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

Thank You So Much